OpenStudy (mckenzieandjesus):

Find the value of x. Round your answer to the nearest tenth. mc056-1.jpg

OpenStudy (mckenzieandjesus):

tan x degrees= 3/5

OpenStudy (mckenzieandjesus):

someone please help ;/

OpenStudy (anonymous):

No calculator?

OpenStudy (mckenzieandjesus):

Yes i have one

OpenStudy (mckenzieandjesus):

I think the answer is 31.0 but not positive

OpenStudy (anonymous):

So hit \[\tan^{-1} (3/5)\]

OpenStudy (anonymous):

It's probably 0

OpenStudy (mckenzieandjesus):

its 0.01047235832

OpenStudy (mckenzieandjesus):

yes 0

OpenStudy (anonymous):

So u know how to round it now?

OpenStudy (anonymous):

Well you deserve the best answer then :P

OpenStudy (mckenzieandjesus):

on my multiple choice there is no 0 answer

OpenStudy (mckenzieandjesus):

thank you for the medal

OpenStudy (mckenzieandjesus):

these are my choices 10.5 31.0 29.1 59.0

OpenStudy (anonymous):

Hmmm, do u know how to find the answer in degrees?it's 0.01 in rads

OpenStudy (mckenzieandjesus):

no i dont..

OpenStudy (mckenzieandjesus):

it be just 1% i think.. but thats not the answer

OpenStudy (anonymous):

OK u have the proptionality given by: Pi stands for 180 ,then one rad stands for 180/3.14 Therefore 0.01 rad= 0.01* (180/3.14) Try it

OpenStudy (mckenzieandjesus):

0.5732484076

OpenStudy (anonymous):

I think u meant 5.7

OpenStudy (mckenzieandjesus):

oh yeah

OpenStudy (anonymous):

But still wrong I see. I think we did nothing wrong. Can urepeat the calculations. Notice that answer A is double what we got

OpenStudy (anonymous):

I can give u a hint mckenzie, the answer is A lol. Just redo the steps one more time please

OpenStudy (mckenzieandjesus):

i got 31.0 from someone else but idk if its right

OpenStudy (anonymous):

The verification is soeasy, just try tan 31 in the calclator

OpenStudy (mckenzieandjesus):

are u sure its A?

OpenStudy (anonymous):

Try both :P

OpenStudy (mckenzieandjesus):

it comes out to 6 something

OpenStudy (anonymous):

3/5 is around .6

OpenStudy (mckenzieandjesus):

3/ is 0.6

OpenStudy (anonymous):

Your answer is about .6 then

OpenStudy (mckenzieandjesus):

3/5

OpenStudy (mckenzieandjesus):

which can't be right theres no 6 answer

OpenStudy (anonymous):

No I mean the original equation, tan(x)=0.6

OpenStudy (anonymous):

So what ever is x, tanx must be 0.6

OpenStudy (mckenzieandjesus):

so are u sure the answer is a cuz i saw that its 31.0 somewhere else

OpenStudy (mckenzieandjesus):

im so confused lol

OpenStudy (mckenzieandjesus):

if 3/5 =0.6 then the answer is 31.0 cuz tan 31.0 equals 0.6?

OpenStudy (anonymous):

I'm sorry for confusion :/ I brought my calculator. The answer is actually 31 lol

OpenStudy (mckenzieandjesus):

thats what i thought

OpenStudy (mckenzieandjesus):

thanku

OpenStudy (anonymous):

I didn't think it would take that much time, sorry for cofusion. I repeat it' just calculator thing Good luck